OpenStudy (anonymous):

Solve x+y=1 and y=2/3x-4 algebrically

OpenStudy (anonymous):

is this a system of equations?

OpenStudy (anonymous):

and is the 2/3 a fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, solve for y in y=2/3x-4

OpenStudy (anonymous):

to still get y=2/3x-4

OpenStudy (anonymous):

then substitute y=2/3x-4 for x+y=1

OpenStudy (anonymous):

x+2/3x-4=1

OpenStudy (anonymous):

solve that to get x=3

OpenStudy (anonymous):

Then what do u do?

OpenStudy (anonymous):

then substitutex=3 in y=2/3x-4

OpenStudy (anonymous):

y=2/3*3-4

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

to get y=-2

OpenStudy (anonymous):

x=3 and y=-2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem :)